# Weak coloring numbers of minor-closed graph classes

### v1

@@ 1. Introduction
$\operatorname{td}(X)=\begin{cases}0&\textrm{if $X$ is the null graph,}\\ \min_{v\in V(X)}\operatorname{td}(X-v)+1&\textrm{if $X$ is connected, and}\\ \max_{i\in[k]}\operatorname{td}(C_{i})&\textrm{if $X$ consists of components $C_{1},\dots,C_{k}$ and $k>1$.}\end{cases}$
@@
The aforementioned results imply that $\operatorname{tw}(X)-1\leqslant f(X)\leqslant\operatorname{td}(X)-1$. However, $f$ is not tied to any of these parameters. Indeed, neither pathwidth nor treedepth can lower bound $f$. For every positive integer $k$, let $T_{k}$ be a complete ternary tree of vertex-height $k$. Recall that there is a constant depending on $k$ bounding pathwidth of $T_{k}$-minor-free graphs by Robertson-Seymour Excluded Tree Minor Theorem [19]. Also it is easy to show that $\operatorname{wcol}_{r}(G)\leqslant(\operatorname{pw}(G)+1)(2r+1)$ for every graph $G$. Thus, $f(T_{k})\leqslant 1$ while $\operatorname{pw}(T_{k})=k$ and $\operatorname{td}(T_{k})=k+1$. Next, we argue that neither treewidth nor pathwidth can upper-bound $f$. For every positive integer $k$, let $L_{k}$ be a ladder with $k$ rungs. There is a graph $G_{r,t}$ (constructed in [8]) such that $\operatorname{wcol}_{r}(G_{r,t})=\Omega(r^{t})$, and if $k=\Omega(\log t)$, then $G_{r,t}$ excludes $L_{k}$ as a minor. Therefore, $f(L_{k})=2^{\Omega(k)}$, and $\operatorname{tw}(L_{k})\leqslant\operatorname{pw}(L_{k})\leqslant 2$.
@@
To prove Theorems 1 and 4, we prove that the value of $f$ is tied with the maximum $t$ such that $X$ is a subgraph of $G_{r,t}$ (as in [8]). In other words, we prove that $G_{r,t}$ is the obstruction for the growth of weak coloring numbers. More precisely, we introduce a slightly modified version of $2$-treedepth, which we call rooted $2$-treedepth and denote by $\operatorname{rtd}_{2}(\cdot)$. Later, we show that for all graphs $X$ with at least one edge, $\operatorname{rtd}_{2}(X)$ the minimum $t$ such that there exists $r$ such that $X$ is a subgraph of $G_{r,t-1}$. See Section 4 for the definition of rooted $2$-treedepth and Lemma 26 for the equivalence.
@@
Table 1. Lower and upper bounds on $\max_{G\in\mathcal{C}}\operatorname{wcol}_{r}(G)$ for some minor-closed graphs classes $\mathcal{C}$. The variables $g,k,s,t$ are fixed positive integers with $s+3\leqslant t\leqslant k$, and $X$ is a fixed nonnull graph. The weak coloring numbers of $K_{s,t}$-minor-free graphs were first studied by van den Heuvel and Wood in [10]. In particular they proved the upper bound $\mathcal{O}(r^{s+1})$ and they conjectured $\mathcal{O}(r^{s})$. Theorem 1 implies that $K_{s,t}$-minor-free graphs have weak coloring numbers in $\mathcal{O}(r^{s-1}\log r)$. The lower bound $\Omega(r^{s-1}\log r)$ follows from the fact that graphs of simple treewidth $s$ are $K_{s,t}$-minor-free and among them there are graphs with weak coloring numbers in $\Omega(r^{s-1}\log r)$, see [13] for further details.
@@ 2.1. The key parameter
Let $G$ be a graph and let $k$ be a nonnegative integer. A separation of order $k$ of $G$ is a pair $(A,B)$ of subgraphs of $G$ such that $A\cup B=G$, $E(A\cap B)=\emptyset$, and $|V(A\cap B)|=k$. We define recursively a new graph parameter called rooted $2$-treedepth, denoted by $\operatorname{rtd}_{2}$, as follows. For every graph $G$,
@@
Another way to understand $\operatorname{rtd}_{2}$ is through “natural” separations of the graph defined by its block decomposition. When $G$ is not connected, $\operatorname{rtd}_{2}(G)$ is realized by $\operatorname{rtd}_{2}(C)$, where $C$ is a component of $G$ for which the value of $\operatorname{rtd}_{2}$ is the greatest. When $G$ consists of a single block, a separation $(A,B)$ of $G$ with $V(A)\neq\emptyset$ and $V(B)-V(A)\neq\emptyset$ of order at most one is such that $V(A)=\{v\}$. Therefore, $\operatorname{rtd}_{2}(G)=\operatorname{rtd}_{2}(B-v)+1$. When $G$ consists of multiple blocks, the minimum of $\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-V(A))+|V(A)\cap V(B)|\}$ is reached for separations where $V(A)\cap V(B)$ consists of exactly one cut-vertex of $G$. The above can be summarized as the following properties. For every graph $G$,
@@
We now recall this construction. Let $d$ be a positive integer, let $B$ and $H$ be two graphs and let $u$ be a vertex of $H$. We define $L_{d}(B,H,u)$ as the graph obtained in the following process. Take a copy of $B$ and $d|V(B)|$ copies of $H$. Label the latter $H_{i,x}$ for each $i\in[d]$ and $x\in V(B)$. Next, for each $x\in V(B)$ identify $x$ and $u$ in each $H_{i,x}$ for $i\in[d]$. See Figure 2.
@@
Therefore, by induction, $\operatorname{rtd}_{2}(G_{r,t})=t+1$. In fact, this construction is universal for graphs of rooted $2$-treedepth at most $t+1$. Namely, for every graph $G$, $\operatorname{rtd}_{2}(G)\leqslant t+1$ if and only if $G$ is isomorphic to a subgraph of $G_{r,t}$ for some nonnegative integer $r$. Since rooted $2$-treedepth is minor-monotone, we deduce that for every graph $X$, $G_{r,\operatorname{rtd}_{2}(X)-2}$ is $X$-minor-free. It follows that there are $X$-minor-free graphs with $r$-th weak coloring numbers in $\Omega(r^{\operatorname{rtd}_{2}(X)-2})$. Hence, Theorem 5 yields that for every nonnegative integer $t$, if a minor-closed class of graphs contains graphs with $r$-th weak coloring in $\omega(r^{t}\log r)$, then it contains $G_{r,t}$ for every nonnegative integer $r$. As a consequence, in the setting of minor-closed graphs classes, our results imply that the family constructed by Grohe et al. is, up to an $\mathcal{O}(r\log r)$ factor, the unique construction of graphs with large weak coloring numbers. Similarly, in the setting of minor-closed graphs classes of bounded treewidth, the family constructed by Grohe et al. is, up to an $\mathcal{O}(\log r)$ factor, the unique construction of graphs with large weak coloring numbers.
@@ 2.3. $\mathcal{F}$-rich models
If $G$ has a model of $H$, then we say that $H$ is a minor of $G$. Let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$. Such a model of $H$ is said to be $\mathcal{F}$-rich if we have the additional following property.
@@ 2.4. Plan of the proof
We now present the main ideas behind the proofs of Theorem 5 and Theorem 6. In order to prove the theorems, we strengthen the statement and instead of considering graphs with no models of $X$, we consider graphs with no $\mathcal{F}$-rich models of $X$, given a family $\mathcal{F}$ of nonnull connected subgraphs of $G$. This turns out to be very helpful in keeping induction invariant if we carefully choose $\mathcal{F}$. For some choices of $\mathcal{F}$, excluding an $\mathcal{F}$-rich model of a graph may be a local property. Therefore, we keep a global property that $G$ excludes $K_{k}$ as a minor, where $k=|V(X)|$. The locality of the property suggests that the bound on weak coloring numbers should be local in some sense as well. To this end, instead of bounding $\operatorname{wcol}_{r}(G)$, we bound $\operatorname{wcol}_{r}(G,S)$, where $S$ is a hitting set for $\mathcal{F}$, that is $V(F)\cap S\neq\emptyset$ for every $F\in\mathcal{F}$. A result in this spirit, that corresponds to excluding every $\mathcal{F}$-rich model of an edgeless graph $X$, is already present in [2]. We restate this result below with adjusted notations.
@@
There exists a function $\delta$ such that for all positive integers $k,d$, for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if there are no $d$ pairwise vertex-disjoint subgraphs in $\mathcal{F}$, then there exists $S\subseteq V(G)$ such that
@@
Let $k$ and $t$ be positive integers with $t\geqslant 2$. Let $X$ be a graph with $\operatorname{rtd}_{2}(X)\leqslant t$. There exists an integer $c(t,X,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $X$, then there exists $S\subseteq V(G)$ such that
@@
Figure 4. On the left-hand side, we depict an $\mathcal{F}$-rich model of $X$, where $X$ is a cycle on $8$ vertices and $\mathcal{F}$ is the family of all nonnull connected subgraphs of $G-\{u\}$ containing a neighbor of $u$ in $G$. On the right-hand side, we show how to construct, given an $\mathcal{F}$-rich model of $Y$, a model of $K_{1}\oplus X$.
@@ 3. Preliminaries
Let $G$ be a graph. For $X,Y\subseteq V(G)$, an $X$-$Y$ path is a path in $G$ that is either a one-vertex path with the vertex in $X\cap Y$ or a path with one endpoint in $X$ and the other endpoint in $Y$ such that no internal vertices are in $X\cup Y$. When $u,v\in V(G)$, instead of $\{u\}$-$\{v\}$ path we write $u$-$v$ path for short. The length of a path $P$ is the number of edges of $P$. A path $P$ is a geodesic in $G$ if it is a shortest path between its endpoints in $G$. The distance between two vertices $u$ and $v$ in $G$, denoted by $\mathrm{dist}_{G}(u,v)$, is the length of a $u$-$v$ geodesic in $G$ when it exists, and +$\infty$+ ∞ otherwise. Let $u$ be a vertex of $G$. The neighborhood of $u$ in $G$, denoted by $N_{G}(u)$, is the set $\{v\in V(G)\mid uv\in E(G)\}$. For every set of vertices $X$ of $G$, let $N_{G}(X)=\bigcup_{u\in X}N_{G}(u)-X$. For every positive integer $r$, we denote by $N_{G}^{r}[u]=\{v\in V(G)\mid\mathrm{dist}_{G}(u,v)\leqslant r\}$. We omit $G$ in the subscripts when it is clear from the context. The following lemma is folklore, see e.g. [2, Lemma 23] for a proof.
@@
This inequality is witnessed by an ordering of $S\cup V(Q_{1})\cup\cdots V(Q_{\ell})$ obtained from the ordering $\sigma$ of $S$ witnessing $\operatorname{wcol}_{r}(G,S)$ by putting vertices from $V(Q_{1})\cup\cdots V(Q_{\ell})-S$ arbitrarily.
@@
There exists a function $\delta$ such that for all positive integers $k,d$, for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, either there are $d$ pairwise vertex-disjoint subgraphs in $\mathcal{F}$, or there exists $S\subseteq V(G)$ such that
@@
There exists a function $\gamma$ such that for all positive integers $k,d$, for every $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$ either
@@
    there exists $A\subseteq V(G)$ such that $|A|\leqslant(d-1)\gamma(k)$, and a subgraph $X$ of $G$, where $X$ is the union of at most $(d-1)^{2}\gamma(k)$ geodesics in $G-A$, and for every $F\in\mathcal{F}$ we have $V(F)\cap(V(X)\cup A)\neq\emptyset$.
@@
Let $\delta(k,d)=12(d-1)^{2}\gamma(k)$ where $\gamma$ is the function from Lemma 18. Let $G$ be a $K_{k}$-minor-free graph and let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$. Suppose that there are no $d$ pairwise disjoint members of $\mathcal{F}$, and hence, Lemma 18.(2) holds, yielding $A\subseteq V(G)$ and a subgraph $X$ of $G$ such that $|A|\leqslant(d-1)\gamma(k)$ and $X$ is the union of at most $(d-1)^{2}\gamma(k)$ geodesics in $G-A$. Note that $G[A\cup V(X)]$ has at most $|A|+(d-1)^{2}\gamma(k)$ components. Let $Q_{1},\dots,Q_{\ell}$ be a family of at most $(d-1)\gamma(k)+(d-1)^{2}\gamma(k)-1$ geodesics in $G$ such that the set $S=A\cup V(X)\cup\bigcup_{i\in[\ell]}V(Q_{i})$ induces a connected subgraph in $G$. In particular, $\ell\leqslant 2(d-1)^{2}\gamma(k)$. By Observations 13 and 16, for every positive integer $r$,
@@
It is easy to derive an upper bound on weak coloring numbers of trees. It suffices to root a given tree and order the vertices in an elimination order. Namely, an elimination ordering of a tree $T$ rooted in $s\in V(T)$ is an ordering $(x_{1},\dots,x_{|V(T)|})$ of $V(T)$ such that $x_{1}=s$ and for every $i\in\{2,\dots,|V(T)|\}$, $N(x_{i})\cap\{x_{j}\mid j\in[i-1]\}=\{y\}$ where $y$ is the parent of $x$. Note that in such an ordering a vertex weakly reaches only its ancestors. See Figure 5.
@@
where $A_{r,x}$ is the set of ancestors of $x$ in distance at most $r$. Note that $|A_{r,x}|\leqslant r+1$.
Note that this idea can be generalized to elimination orderings of tree decompositions – see [8].

### v2

@@ 1. Introduction
[…]
@@
The aforementioned results imply that $\operatorname{tw}(X)-1\leqslant f(X)\leqslant\operatorname{td}(X)-1$. However, $f$ is not tied to any of these parameters. Indeed, neither pathwidth nor treedepth can lower bound $f$. For every positive integer $k$, let $T_{k}$ be a complete ternary tree of vertex-height $k$. Recall that there is a constant depending on $k$ bounding pathwidth of $T_{k}$-minor-free graphs by Robertson-Seymour Excluded Tree Minor Theorem [19]. Also, it is easy to show [^5] Proceed by induction on $\operatorname{pw}(G)$. We may assume that $G$ is connected. If $\operatorname{pw}(G)=0$, then $G$ has no edge and so $\operatorname{wcol}_{r}(G)\leqslant 1$. If $\operatorname{pw}(G)>0$, let $Q$ be a shortest path from the first bag to the last bag of an optimal path decomposition of $G$. Then $\operatorname{pw}(G-V(Q))<\operatorname{pw}(G)$ and so by induction $\operatorname{wcol}_{r}(G-V(Q))\leqslant 1+(\operatorname{pw}(G)-1)(2r+1)$. Let $\sigma_{0}$ be an ordering of $V(G)$ witnessing this fact. Now, let $\sigma$ be an ordering of $V(G)$ extending $\sigma_{0}$ such that the vertices in $V(Q)$ appear first. By Lemma 9, it follows that $\sigma$ witnesses $\operatorname{wcol}_{r}(G)\leqslant 1+(\operatorname{pw}(G)-1)(2r+1)+(2r+1)=1+\operatorname{pw}(G)(2r+1)$. that $\operatorname{wcol}_{r}(G)\leqslant 1+\operatorname{pw}(G)(2r+1)$ for every graph $G$. Thus, $f(T_{k})\leqslant 1$ while $\operatorname{pw}(T_{k})=k$ and $\operatorname{td}(T_{k})=k+1$. Next, we argue that neither treewidth nor pathwidth can upper-bound $f$. For every positive integer $k$, let $L_{k}$ be a ladder with $k$ rungs. There is a graph $G_{r,t}$ (constructed in [8]) such that $\operatorname{wcol}_{r}(G_{r,t})=\Omega(r^{t})$, and if $k=\Omega(\log t)$, then $G_{r,t}$ excludes $L_{k}$ as a minor. Therefore, $f(L_{k})=2^{\Omega(k)}$, and $\operatorname{tw}(L_{k})\leqslant\operatorname{pw}(L_{k})\leqslant 2$.
@@
To prove Theorems 1 and 4, we prove that the value of $f$ is tied with the maximum $t$ such that $X$ is a subgraph of $G_{r,t}$ (as in [8]). In other words, we prove that $G_{r,t}$ is the obstruction for the growth of weak coloring numbers. More precisely, we introduce a slightly modified version of $2$-treedepth, which we call rooted $2$-treedepth and denote by $\operatorname{rtd}_{2}(\cdot)$. Later, we show that for all graphs $X$ with at least one edge, $\operatorname{rtd}_{2}(X)$ is the minimum $t$ such that there exists $r$ such that $X$ is a subgraph of $G_{r,t-1}$. See Section 4 for the definition of rooted $2$-treedepth and Lemma 26 for the equivalence.
@@
Table 1. Lower and upper bounds on $\max_{G\in\mathcal{C}}\operatorname{wcol}_{r}(G)$ for some minor-closed graphs classes $\mathcal{C}$. The variables $g,k,s,t$ are fixed positive integers with $s+3\leqslant t\leqslant k$, and $X$ is a fixed nonnull graph. The weak coloring numbers of $K_{s,t}$-minor-free graphs were first studied by van den Heuvel and Wood in [10]. In particular they proved the upper bound $\mathcal{O}(r^{s+1})$ and they conjectured $\mathcal{O}(r^{s})$. Theorem 1 implies that $K_{s,t}$-minor-free graphs have weak coloring numbers in $\mathcal{O}(r^{s}\log r)$. The lower bound $\Omega(r^{s-1}\log r)$ follows from the fact that graphs of simple treewidth $s$ are $K_{s,t}$-minor-free and among them there are graphs with weak coloring numbers in $\Omega(r^{s-1}\log r)$, see [13] for further details.
@@ 2.1. The key parameter
Let $G$ be a graph and let $k$ be a nonnegative integer. A separation of order $k$ of $G$ is a pair $(A,B)$ of subgraphs of $G$ such that [^8] For all graphs $A,B$, let $A\cup B=(V(A)\cup V(B),E(A)\cup E(B))$ and $A\cap B=(V(A)\cap V(B),E(A)\cap E(B))$. $A\cup B=G$, $E(A\cap B)=\emptyset$, and $|V(A\cap B)|=k$. We define recursively a new graph parameter called rooted $2$-treedepth, denoted by $\operatorname{rtd}_{2}$, as follows. For every graph $G$,
@@
Another way to understand $\operatorname{rtd}_{2}$ is through “natural” separations of the graph defined by its block decomposition. When $G$ is not connected, $\operatorname{rtd}_{2}(G)$ is realized by $\operatorname{rtd}_{2}(C)$, where $C$ is a component of $G$ for which the value of $\operatorname{rtd}_{2}$ is the greatest. When $G$ consists of a single block, a separation $(A,B)$ of $G$ with $V(A)\neq\emptyset$ and $V(B)-V(A)\neq\emptyset$ of order at most one is such that $V(A)=\{v\}$ and $B=G$. Therefore, $\operatorname{rtd}_{2}(G)=\operatorname{rtd}_{2}(G-v)+1$. When $G$ consists of multiple blocks, the minimum of $\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-V(A))+|V(A)\cap V(B)|\}$ is reached for separations where $V(A)\cap V(B)$ consists of exactly one cut-vertex of $G$. The above can be summarized as the following properties. For every graph $G$,
@@
We now recall this construction. Let $d$ be a positive integer, let $B$ and $H$ be two graphs, and let $u$ be a vertex of $H$. We define $L_{d}(B,H,u)$ as the graph obtained in the following process. Take a copy of $B$ and $d|V(B)|$ copies of $H$. Label the latter $H_{i,x}$ for each $i\in[d]$ and $x\in V(B)$. Next, for each $x\in V(B)$ identify $x$ and $u$ in each $H_{i,x}$ for $i\in[d]$. See Figure 2.
@@
Therefore, by induction, $\operatorname{rtd}_{2}(G_{r,t})=t+1$. In fact, this construction is universal for graphs of rooted $2$-treedepth at most $t+1$ (see Lemma 26). Namely, for every graph $G$, $\operatorname{rtd}_{2}(G)\leqslant t+1$ if and only if $G$ is isomorphic to a subgraph of $G_{r,t}$ for some nonnegative integer $r$. Since rooted $2$-treedepth is minor-monotone, we deduce that for every graph $X$, $G_{r,\operatorname{rtd}_{2}(X)-2}$ is $X$-minor-free. It follows that there are $X$-minor-free graphs with $r$-th weak coloring numbers in $\Omega(r^{\operatorname{rtd}_{2}(X)-2})$. Hence, Theorem 5 yields that for every nonnegative integer $t$, if a minor-closed class of graphs contains graphs with $r$-th weak coloring in $\omega(r^{t}\log r)$, then it contains $G_{r,t}$ for every nonnegative integer $r$. As a consequence, in the setting of minor-closed graphs classes, our results imply that the family constructed by Grohe et al. is, up to an $\mathcal{O}(r\log r)$ factor, the unique construction of graphs with large weak coloring numbers. Similarly, in the setting of minor-closed graphs classes of bounded treewidth, the family constructed by Grohe et al. is, up to an $\mathcal{O}(\log r)$ factor, the unique construction of graphs with large weak coloring numbers.
@@ 2.3. $\mathcal{F}$-rich models
If $G$ has a model of $H$, then we say that $H$ is a minor of $G$. Let $\mathcal{F}$ be a family of connected subgraphs of $G$. Such a model of $H$ is said to be $\mathcal{F}$-rich if we have the additional following property.
@@ 2.4. Plan of the proof
We now present the main ideas behind the proofs of Theorem 5 and Theorem 6. In order to prove the theorems, we strengthen the statement and instead of considering graphs with no models of $X$, we consider graphs with no $\mathcal{F}$-rich models of $X$, given a family $\mathcal{F}$ of connected subgraphs of $G$. This turns out to be very helpful in keeping induction invariant if we carefully choose $\mathcal{F}$. For some choices of $\mathcal{F}$, excluding an $\mathcal{F}$-rich model of a graph may be a local property. Therefore, we keep a global property that $G$ excludes $K_{k}$ as a minor, where $k=|V(X)|$. The locality of the property suggests that the bound on weak coloring numbers should be local in some sense as well. To this end, instead of bounding $\operatorname{wcol}_{r}(G)$, we bound $\operatorname{wcol}_{r}(G,S)$, where $S$ is a hitting set for $\mathcal{F}$, that is $V(F)\cap S\neq\emptyset$ for every $F\in\mathcal{F}$. A result in this spirit, that corresponds to excluding every $\mathcal{F}$-rich model of an edgeless graph $X$, is already present in [2]. We restate this result below with adjusted notations (see Lemma 18 for the original statement).
@@
There exists a function $\delta$ such that for all positive integers $k,d$, for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if there are no $d$ pairwise vertex-disjoint subgraphs in $\mathcal{F}$, then there exists $S\subseteq V(G)$ such that
@@
Let $k$ and $t$ be positive integers with $t\geqslant 2$. Let $X$ be a graph with $\operatorname{rtd}_{2}(X)\leqslant t$. There exists an integer $c(t,X,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $X$, then there exists $S\subseteq V(G)$ such that
@@
Figure 4. On the left-hand side, we depict an $\mathcal{F}$-rich model of $X$, where $X$ is a cycle on $8$ vertices and $\mathcal{F}$ is the family of all connected subgraphs of $G-\{u\}$ containing a neighbor of $u$ in $G$. On the right-hand side, we show how to construct, given an $\mathcal{F}$-rich model of $X$, a model of $K_{1}\oplus X$.
@@ 3. Preliminaries
Let $G$ be a graph. For $X,Y\subseteq V(G)$, an $X$-$Y$ path is a path in $G$ that is either a one-vertex path with the vertex in $X\cap Y$ or a path with one endpoint in $X$ and the other endpoint in $Y$ such that no internal vertices are in $X\cup Y$. When $u,v\in V(G)$, instead of $\{u\}$-$\{v\}$ path we write $u$-$v$ path for short. The length of a path $P$ is the number of edges of $P$. A path $P$ is a geodesic in $G$ if it is a shortest path between its endpoints in $G$. The distance between two vertices $u$ and $v$ in $G$, denoted by $\mathrm{dist}_{G}(u,v)$, is the length of a $u$-$v$ geodesic in $G$ when it exists, and +$\infty$+ ∞ otherwise. Let $u$ be a vertex of $G$. The neighborhood of $u$ in $G$, denoted by $N_{G}(u)$, is the set $\{v\in V(G)\mid uv\in E(G)\}$. For every set of vertices $X$ of $G$, let $N_{G}(X)=\bigcup_{u\in X}N_{G}(u)-X$. For every positive integer $r$, we denote by $N_{G}^{r}[u]=\{v\in V(G)\mid\mathrm{dist}_{G}(u,v)\leqslant r\}$. The following lemma is folklore, see e.g. [2, Lemma 23] for a proof.
@@
This inequality is witnessed by an ordering of $S\cup V(Q_{1})\cup\cdots V(Q_{\ell})$ obtained from the ordering $\sigma$ of $S$ witnessing $\operatorname{wcol}_{r}(G,S)$ by putting vertices from $(V(Q_{1})\cup\cdots\cup V(Q_{\ell}))-S$ arbitrarily.
@@
There exists a function $\delta$ such that for all positive integers $k,d$, for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, either there are $d$ pairwise vertex-disjoint subgraphs in $\mathcal{F}$, or there exists $S\subseteq V(G)$ such that
@@
There exists a function $\gamma$ such that for all positive integers $k,d$, for every $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$ either
@@
    there exists $A\subseteq V(G)$ with $|A|\leqslant(d-1)\gamma(k)$ and there exists a subgraph $X$ of $G$ which is the union of at most $(d-1)^{2}\gamma(k)$ geodesics in $G-A$, such that for every $F\in\mathcal{F}$ we have $V(F)\cap(V(X)\cup A)\neq\emptyset$.
@@
Let $\delta(k,d)=12(d-1)^{2}\gamma(k)$ where $\gamma$ is the function from Lemma 18. Let $G$ be a $K_{k}$-minor-free graph and let $\mathcal{F}$ be a family of connected subgraphs of $G$. Suppose that there are no $d$ pairwise disjoint members of $\mathcal{F}$, and hence, Lemma 18.(2) holds, yielding $A\subseteq V(G)$ and a subgraph $X$ of $G$ such that $|A|\leqslant(d-1)\gamma(k)$ and $X$ is the union of at most $(d-1)^{2}\gamma(k)$ geodesics in $G-A$. Note that $G[A\cup V(X)]$ has at most $|A|+(d-1)^{2}\gamma(k)$ components. Let $Q_{1},\dots,Q_{\ell}$ be a family of at most $(d-1)\gamma(k)+(d-1)^{2}\gamma(k)-1$ geodesics in $G$ such that the set $S=A\cup V(X)\cup\bigcup_{i\in[\ell]}V(Q_{i})$ induces a connected subgraph in $G$. In particular, $\ell\leqslant 2(d-1)^{2}\gamma(k)$. By Observations 13 and 16, for every positive integer $r$,
@@
It is easy to derive an upper bound on weak coloring numbers of trees. It suffices to root a given tree and order the vertices in an elimination order. Namely, an elimination ordering of a tree $T$ rooted in $s\in V(T)$ is an ordering $(x_{1},\dots,x_{|V(T)|})$ of $V(T)$ such that $x_{1}=s$ and for every $i\in\{2,\dots,|V(T)|\}$, $N(x_{i})\cap\{x_{j}\mid j\in[i-1]\}=\{y\}$ where $y$ is the parent of $x_{i}$. Note that in such an ordering a vertex weakly reaches only its ancestors. See Figure 5.
@@
where $A_{r,x}$ is the set of ancestors of $x$, including $x$ itself, in distance at most $r$. Note that $|A_{r,x}|\leqslant r+1$.
This idea can be generalized to elimination orderings of tree decompositions – see [8].